# Utilize the Flow before Stepping into the Same River Twice: Certainty Represented Knowledge Flow for Refusal-Aware Instruction Tuning

### v2

@@ Abstract
Refusal-Aware Instruction Tuning (RAIT) enables Large Language Models (LLMs) to refuse to answer unknown questions. By modifying responses of unknown questions in the training data to refusal responses such as “I don’t know”, RAIT enhances the reliability of LLMs and reduces their hallucination. Generally, RAIT modifies training samples based on the correctness of the initial LLM’s response. However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal. In this paper, we explore two primary causes of over-refusal: Static conflict occurs when similar samples within the LLM’s feature space receive differing supervision signals (original vs. modified “I don’t know”). Dynamic conflict, on the other hand, emerges as the LLM’s knowledge evolves during SFT, allowing it to answer questions that were previously unanswerable. Yet, these now-answerable training samples still retain the original “I don’t know” supervision signals based on the initial LLM state, resulting in inconsistencies. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal. To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Tuning (CRaFT). CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts. Second, we implement preliminary rehearsal training to characterize changes in the LLM’s knowledge state, which helps mitigate dynamic conflicts during the fine-tuning process. We conducted extensive experiments on open-ended question answering and multiple-choice question task. Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process. Source code and training data will be released at Github.
@@ 1 Introduction
Recently, Large Language Model (LLM) technology has made significant progress, becoming an important milestone towards AGI (Achiam et al. 2023; Dubey et al. 2024; Yang et al. 2024). However, current LLMs often output fabricated information, which is referred to as hallucinations (Ji et al. 2023). This phenomenon severely limits the usefulness and reliability of LLMs. An important reason for the occurrence of hallucinations is that when exposed to questions beyond their internal knowledge (i.e., unknown questions), LLMs may forcefully generate responses (Kang et al. 2024). Ideally, the reliable LLM should actively refuse to answer questions it doesn’t know to avoid incorrect responses (Wen et al. ; Li et al. 2024b). Recent studies have shown (Yang et al. 2023; Zhang et al. 2024; Xu et al. 2024; Cheng et al. 2024) that Refusal-Aware Instruction Tuning (RAIT) can enable LLMs to refuse answering questions beyond their knowledge.
@@
Generally, the RAIT process cab be described as follows: the initial LLM is prompted to answer all questions in the train set $D_{\text{src}}$. Based on the response correctness, the samples are categorized into two groups. Samples with correct responses are considered as known knowledge of the LLM. Their answers will not be modified, referred to as vanilla samples $D_{\text{van}}$. Conversely, samples with incorrect responses are considered as unknown knowledge of LLM. Their original answers are replaced with refusal responses like “I don’t know”. These samples are called IdK samples $D_{\text{idk}}$. The constructed RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$ is used to fine-tune the initial LLM, improving its ability to refuse to answer unknown questions. For simplicity, we refer to this original form of the RAIT method as Cor-RAIT.
However, (Cheng et al. 2024) reveals that the Cor-RAIT causes the fine-tuned LLM to refuse some questions that could have been answered correctly. Experiments demonstrate significant drop of LLM’s accuracy after Cor-RAIT: on the TriviaQA dataset (Joshi et al. 2017), it decreases from 45.05% to 28.57%. Similarly, on the Natural Questions dataset (Kwiatkowski et al. 2019), accuracy fell from 24.65% to 15.93%. We refer to this phenomenon as over-refusal, as shown in Figure 1.
@@
In addressing the over-refusal brought by Cor-RAIT, we identified two primary causes as shown in Figure 2. (1) Static Conflict: In the LLM feature space, two closely located samples might be assigned to $D_{\text{van}}$ and $D_{\text{idk}}$ under the Cor-RAIT framework. As illustrated by t-SNE in Figure 3(a), significant intersections exist between $D_{\text{van}}$ and $D_{\text{idk}}$, complicating their differentiation. These similar samples provide conflict supervision during training, impairing the LLM’s ability to distinguish between known and unknown questions, resulting in over-refusal. (2) Dynamic Conflict: This arises from overlooking the dynamic shifts in LLM’s knowledge state during training. Research (Ren et al. 2024; Ren and Sutherland 2024) shows that the knowledge state of LLMs changes during Supervised Fine-Tuning (SFT), with questions potentially shifting from unknown to known and vice versa. This phenomenon is reminiscent of Heraclitus’ saying, “no man ever steps in the same river twice.” However, current methods use static RAIT data reflecting the initial LLM’s knowledge state throughout SFT, which ignores these changes. This oversight leads to conflicts between the RAIT data and the LLM’s evolving knowledge, resulting in nefficient training and over-refusal.
@@
To address the two problems above, we propose Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT). Our approach consists of two stages. Stage 1: Query the knowledge state and flow of LLM. First, the knowledge state of the initial LLM is probed. Compared to Cor-RAIT, we introduce response certainty on top of the LLM’s response correctness. The joint use of these two indicator variables effectively alleviates the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$. Further, to capture the dynamic changes in LLM’s knowledge state during training, we introduce a rehearsal training mechanism. This mechanism fine-tunes the initial LLM using data samples highly consistent with LLM’s internal knowledge. By not introducing any new knowledge during training(Ren et al. 2024; Kang et al. 2024), this method enables us to observe the LLM’s natural internal adjustments to its knowledge state. The difference in knowledge states between the fine-tuned and the initial LLMs helps characterize the knowledge flow within the LLM during training, aiding in the identification and resolution of dynamic conflicts. Stage 2: Refusal-aware instructions construction and tuning. Taking into account both the static knowledge state and dynamic knowledge flow, we filter out vanilla samples and IdK samples from the original RAIT data, therefore effectively reduce both of the static and dynamic conflicts. Finally, we fine-tune the initial LLM using the acquired RATI data, achieving robust overall performance.
@@
In conducting our experimental analysis, we sought a well-founded metric within current research. However, existing methods have notable limitations. For instance, (Feng et al. 2024) proposed using multiple evaluation metrics simultaneously, which is challenging to manage as it is difficult for a model to optimize across multiple criteria at once. Additionally, we provided counterexamples to demonstrate the inadequacy of metrics proposed in other studies. Consequently, we examined these shortcomings and introduced a singular and comprehensive metric: Truthful Helpfulness Score(THS).
@@
  • To address these static and dynamic conflicts, we introduce CRaFT: during the construction of RAIT data, instead of solely depending on correctness, we additionally incorporate certainty to reduce static conflicts; through rehearsal training, we capture the trends of the LLM’s knowledge flow, which helps effectively mitigate dynamic conflicts. Extensive experiments show that CRaFT can effectively alleviate over-refusal and enhance the overall performance of the LLMs.
  • Given the shortcomings of existing refusal-aware metrics, we developed the Truthful Helpfulness Score (THS). This metric balances the reliability and helpfulness of LLMs, enabling a thorough and precise evaluation of their overall performance.
@@ 2.1 Mitigating Hallucinations of LLMs
Researchers have developed several methods to mitigate hallucinations of LLMs, including data augmentation(Neeman et al. 2022), improving decoding strategies(Holtzman et al. 2019; Chuang et al. 2023), integrating external knowledge (Karpukhin et al. 2020), knowledge editing (Zhang, Yu, and Feng 2024; Li et al. 2024a) and honesty alignment(Zhang et al. 2024; Xu et al. 2024). Unlike the previous approach of correcting model answers, honesty alignment encourages models to say “I don’t know” when answering unknown questions. Ways to achieve honesty alignment include refusal-aware instruction tuning (RAIT) and reinforcement learning-based honesty alignment (RLHA).
@@ 2.2 Refusal-Aware Instruction Tuning
RAIT is the supervised technique that improves LLMs’ responses by training LLMs to directly respond with “I don’t know” to unknown questions. R-Tuning (Zhang et al. 2024) identifies these unknown questions by having the LLM answer each question in the training set a single time and then verifying whether the LLM’s responses match the label. In (Yang et al. 2023), the LLM is prompted to answer the same question multiple times. The target answer is then modified based on the correctness ratio of the responses. In (Wan et al. 2024), a knowledge-based verification mechanism ensures the model’s knowledge to be consistent with external trusted sources, increasing refusal accuracy and preventing the spread of false information.
2.3 Reinforcement Learning-based Honesty Alignment
RLHA improves model refusal ability by reinforcement learning with well-designed reward models. (Xu et al. 2024) proposed a novel alignment framework called RLKF, which used knowledge feedback to dynamically determine the knowledge boundary of LLM and train reliable reward models to encourage refusal responses. (Cheng et al. 2024) introduces a refusal function that learns whether the model should refuse to answer by evaluating the extent to which the input aligns with human honesty.
@@ 3.1 Refusal-Aware Instruction Tuning
Given the initial LLM $\mathcal{M}_{0}$ and the instruction dataset $D_{\text{src}}$ consisting of question and answer pairs $x=(q,a)$, we select and modify data from $D_{\text{src}}$ to construct the data $D_{\text{rait}}$, composed of pairs $(q,a_{\text{rait}})$. $D_{\text{rait}}$ is then used to perform SFT on $\mathcal{M}_{0}$, resulting in the new LLM with the capability to decline unknown questions. This process is referred to as Refusal-Aware Instruction Tuning (RAIT). Existing studies (Zhang et al. 2024; Yang et al. 2023; Cheng et al. 2024) use $\mathcal{M}_{0}$ infer questions from $D_{\text{src}}$ and assess their correctness, denoted as $\mu$. As illustrated in Figure 4(a), the correctness threshold $\tau_{\mu}$ is defined firstly. Then, if $\mu<\tau_{\mu}$, the answer for the question is transformed into a new target answer, “I don’t know” and assigned to the subset $D_{\text{idk}}$, comprising what are termed as IdK samples. Conversely, if $\mu\geq\tau_{\mu}$, the original answer remains unchanged and is assigned to the subset $D_{\text{van}}$, which includes the so-called vanilla samples. The resulting RAIT data $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$. This Correctness-based RAIT is referred to as Cor-RAIT.
@@ 3.2 Static Conflicts in Cor-RAIT
During the RAIT process, LLMs learn to reject unknown samples by supervisions from $D_{\text{idk}}$. Our insight is that if the Cor-RAIT dataset contains vanilla and IdK samples that are closely positioned in the LLM’s representation space, the trained LLM may mistakenly classify similar vanilla samples (known questions) as IdK samples (unknown questions), causing over-refusals. To verify this, we analyzed the sample distributions of $D_{\text{van}}$ and $D_{\text{idk}}$. We extract latent representation of each question from the last hidden layer of the LLM. Then, t-SNE is adopted to visualize sample representations. Figure 3(a) displays the distributions of samples from the test split of MMLU dataset (Hendrycks et al. 2020) in the LLaMA-3-8B-instruct (Dubey et al. 2024) representation space, where IdK and vanilla samples have significant intersections.  [^1] We combined additional datasets and LLMs; see Appendix A.1 for similar observations.
@@
The above analysis reveals that Cor-RAIT generates numerous similar sample pairs between $D_{\text{van}}$ and $D_{\text{idk}}$, resulting in conflicting supervision signals which leads to over-refusal. We term this “static conflict” to distinguish it from another conflict type discussed later.
@@ 3.3 Certainty Mitigates the Static Conflicts
We conducted a theoretical analysis 333detailed proof in Appendix A.2 establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$. This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively. To mitigate this, we propose incorporating a robust indicator variable aligned with correctness to select and construct the RAIT data. This variable should ensure that similar samples share comparable values, reducing the above misclassification. We suggest adopting the certainty of the LLM’s response as the indicator variable. Our theoretical analysis shows that certainty meets the above requirements. [^4] Section 4.2 discusses various methods for representing LLM response certainty. Appendix A.3 uses entropy as a measure, but our findings extend to other methods as well.
@@ 3.4 Knowledge Flow and Dynamic Conflict
Research (Ren et al. 2024; Gekhman et al. 2024; Ren and Sutherland 2024) reveals that the knowledge state of LLMs evolves during the SFT process. The phenomenon, which we refer to as “knowledge flow”, can cause previously incorrectly answered questions to become correct ones during SFT. Despite this dynamic evolutions, the target answer of the training data remains static during the SFT process, which reflects the knowledge state of the initial LLM but ignores subsequent changes. We term it as “dynamic conflict” of RAIT data, which significantly contributes to the over-refusal in Cor-RAIT.
We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2. Rehearsal training is designed to capture the LLM’s natural knowledge flow during SFT. Experiments on the MMLU dataset (Hendrycks et al. 2020) and LLaMA-3-8B-instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict. Additional experimental results are provided in Appendix A.5.

### v3

@@ Abstract
Refusal-Aware Instruction Tuning (RAIT) enables Large Language Models (LLMs) to refuse to answer unknown questions. By modifying responses of unknown questions in the training data to refusal responses such as “I don’t know”, RAIT enhances the reliability of LLMs and reduces their hallucination. Generally, RAIT modifies training samples based on the correctness of the initial LLM’s response. However, this crude approach can cause LLMs to excessively refuse answering questions they could have correctly answered, the problem we call over-refusal. In this paper, we explore two primary causes of over-refusal: Static conflict occurs when similar samples within the LLM’s feature space receive differing supervision signals (original vs. modified “I don’t know”). Dynamic conflict arises as the LLM’s evolving knowledge during SFT enables it to answer previously unanswerable questions, but the now-answerable training samples still retain the original “I don’t know” supervision signals from the initial LLM state, leading to inconsistencies. These conflicts cause the trained LLM to misclassify known questions as unknown, resulting in over-refusal. To address this issue, we introduce Certainty Represented Knowledge Flow for Refusal-Aware Instructions Tuning (CRaFT). CRaFT centers on two main contributions: First, we additionally incorporate response certainty to selectively filter and modify data, reducing static conflicts. Second, we implement preliminary rehearsal training to characterize changes in the LLM’s knowledge state, which helps mitigate dynamic conflicts during the fine-tuning process. We conducted extensive experiments on open-ended question answering and multiple-choice question task. Experiment results show that CRaFT can improve LLM’s overall performance during the RAIT process.
Code & Datahttps://github.com/opendatalab/CRaFT
@@ 1 Introduction
Recently, Large Language Model (LLM) technology has made significant progress, becoming an important milestone towards AGI (Achiam et al. 2023; Dubey et al. 2024; Touvron et al. 2023; Yang et al. 2024). However, current LLMs often output fabricated information, which is referred to as hallucinations (Ji et al. 2023). This phenomenon severely limits the usefulness and reliability of LLMs. An important reason for the occurrence of hallucinations is that when exposed to questions beyond their internal knowledge (i.e., unknown questions), LLMs may forcefully generate responses (Kang et al. 2024). Ideally, the reliable LLM should actively refuse to answer questions it doesn’t know to avoid incorrect responses (Wen et al. 2024; Li et al. 2024b). Recent studies have shown (Yang et al. 2023; Zhang et al. 2024; Xu et al. 2024a; Cheng et al. 2024; Xu et al. 2024a; Bai et al. 2023; Cheng et al. 2024) that Refusal-Aware Instruction Tuning (RAIT) can enable LLMs to refuse answering questions beyond their knowledge.
@@
Generally, the RAIT process cab be described as follows: the initial LLM is prompted to answer all questions in the train set $D_{\text{src}}$. Based on the correctness of the responses, the samples are divided into two groups. Correct responses are considered known knowledge and are labeled as vanilla samples $D_{\text{van}}$, whose answers remain unchanged. Incorrect responses are considered unknown knowledge and are replaced with ”I don’t know”, forming the IdK samples $D_{\text{idk}}$. The combined RAIT data, $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$, is used to fine-tune the LLM, enhancing its ability to refuse unknown questions. For simplicity, we refer to this original form of the RAIT method as Cor-RAIT.
@@
However, (Cheng et al. 2024) shows that Cor-RAIT causes the fine-tuned LLM to refuse some questions that could have been answered correctly. Experiments reveal a significant accuracy drop after Cor-RAIT: on the TriviaQA dataset (Joshi et al. 2017), accuracy falls from 45.05% to 28.57%, and on the Natural Questions dataset (Kwiatkowski et al. 2019), it drops from 24.65% to 15.93%. We refer to this phenomenon as over-refusal, as shown in Figure 1.
@@
In addressing the over-refusal brought by Cor-RAIT, we identified two primary causes as shown in Figure 2. (1) Static Conflict: In the LLM representation space, two closely located samples might be assigned to $D_{\text{van}}$ and $D_{\text{idk}}$ under the Cor-RAIT framework. As illustrated by t-SNE in Figure 3(a), significant intersections exist between $D_{\text{van}}$ and $D_{\text{idk}}$, complicating their differentiation. These similar samples provide conflict supervision during training, impairing the LLM’s ability to distinguish between known and unknown questions, resulting in over-refusal. (2) Dynamic Conflict: This arises from overlooking the dynamic shifts in LLM’s knowledge state during training. Research (Ren et al. 2024; Ren and Sutherland 2024; Xu et al. 2024b) shows that the knowledge state of LLMs changes during Supervised Fine-Tuning (SFT), with questions potentially shifting from unknown to known and vice versa. This phenomenon is reminiscent of Heraclitus’ saying, “no man ever steps in the same river twice.” However, current methods use static RAIT data reflecting the initial LLM’s knowledge state throughout SFT, which ignores these changes. This oversight leads to conflicts between the RAIT data and the LLM’s evolving knowledge, resulting inefficient training and over-refusal.
@@
To address the two problems above, we propose Certainty Represented Knowledge Flow for Refusal-Aware Instructions Construction (CRaFT). Our approach consists of two stages. Stage 1: Querying the Knowledge State and Flow of the LLM. First, we probe the initial LLM’s knowledge state. Unlike Cor-RAIT, we incorporate response certainty alongside correctness, effectively alleviating the static conflict between the supervision signals in $D_{\text{van}}$ and $D_{\text{idk}}$. To capture the LLM’s dynamic knowledge changes during training, we introduce a rehearsal training mechanism. This fine-tunes the LLM with data samples that align closely with its internal knowledge, without introducing new knowledge (Ren et al. 2024; Kang et al. 2024). This approach allows us to observe the LLM’s natural knowledge adjustments. The differences between the fine-tuned and initial LLMs reveal the knowledge flow during training, helping to identify and resolve dynamic conflicts. Stage 2: Refusal-aware instructions construction and tuning. By considering both the static knowledge state and dynamic knowledge flow, we filter out vanilla and IdK samples from RAIT data, reducing conflicts. We then fine-tune the initial LLM with the refined data, improving overall performance.
@@
In conducting our experimental analysis, we sought a well-founded metric within current research. Existing methods have notable limitations, either proposing multiple metrics that are hard to optimize simultaneously or relying on inherently flawed metrics, as demonstrated by our counterexamples. Consequently, we examined these shortcomings and introduced a singular and comprehensive metric: Truthful Helpfulness Score(THS).
@@
  • To address static and dynamic conflicts, we introduce CRaFT: it reduces static conflicts by incorporating certainty alongside correctness during RAIT data construction, and mitigates dynamic conflicts through rehearsal training to capture knowledge flow trends. Extensive experiments demonstrate that CRaFT alleviates over-refusal and improves overall LLM performance.
  • We analyze the shortcomings of existing refusal-aware metrics and introduce the Truthful Helpfulness Score (THS), which balances reliability and helpfulness for a comprehensive evaluation of LLM performance.
@@ 2.1 Mitigating Hallucinations of LLMs
Researchers have developed various methods to mitigate LLM hallucinations, including data augmentation (Neeman et al. 2022), improved decoding strategies (Holtzman et al. 2019; Chuang et al. 2023), external knowledge integration (Karpukhin et al. 2020), knowledge editing (Zhang, Yu, and Feng 2024; Li et al. 2024a), and honesty alignment (Zhang et al. 2024; Xu et al. 2024a; Bai et al. 2024). Unlike traditional correction methods, honesty alignment encourages models to say “I don’t know” for unknown questions.
@@ 2.2 Refusal-Aware Instruction Tuning
RAIT is the supervised technique that improves LLMs’ responses by training LLMs to directly respond with “I don’t know” to unknown questions. R-Tuning (Zhang et al. 2024) identifies these questions by having the LLM answer each once and verifying response accuracy. In (Yang et al. 2023), the LLM answers the same question multiple times, with the target answer adjusted based on the correctness ratio. (Wan et al. 2024) uses a knowledge-based verification mechanism to ensure consistency with trusted external sources, enhancing refusal accuracy and preventing misinformation.
@@ 3.1 Refusal-Aware Instruction Tuning
Given the initial LLM $\mathcal{M}_{0}$ and the instruction dataset $D_{\text{src}}$ of question-answer pairs $x=(q,a)$, we modify $D_{\text{src}}$ to construct $D_{\text{rait}}$, consisting of pairs $(q,a_{\text{rait}})$. $D_{\text{rait}}$ is then used for SFT on $\mathcal{M}_{0}$, resulting in a new LLM capable of declining unknown questions, a process called Refusal-Aware Instruction Tuning (RAIT). Existing studies (Zhang et al. 2024; Yang et al. 2023; Cheng et al. 2024) use $\mathcal{M}_{0}$ to infer and assess the correctness of questions in $D_{\text{src}}$, denoted as $\mu$. As shown in Figure 4(a), the correctness threshold $\tau_{\mu}$ is first defined. If $\mu<\tau_{\mu}$, the answer is changed to ”I don’t know” and assigned to the IdK subset $D_{\text{idk}}$. If $\mu\geq\tau_{\mu}$, the original answer remains, and the pair is assigned to the vanilla subset $D_{\text{van}}$. The resulting RAIT data is $D_{\text{rait}}=D_{\text{van}}\cup D_{\text{idk}}$, and this correctness-based RAIT is called Cor-RAIT.
@@ 3.2 Static Conflicts in Cor-RAIT
During the Cor-RAIT process, LLMs learn to reject unknown samples by supervisions from $D_{\text{idk}}$. Our insight is that if the Cor-RAIT dataset contains vanilla and IdK samples that are closely positioned in the LLM’s representation space, the trained LLM may mistakenly classify similar vanilla samples as IdK samples, causing over-refusals. To verify this, we analyzed the sample distributions of $D_{\text{van}}$ and $D_{\text{idk}}$. We extract latent representation of each question from the last hidden layer of the LLM. Then, t-SNE is adopted to visualize sample representations. Figure 3(a) displays the distributions of samples from the test split of MMLU dataset (Hendrycks et al. 2020) in the LLaMA-3-8B-instruct (Dubey et al. 2024) representation space, where IdK and vanilla samples have significant intersections.  [^1] More experiments involving additional datasets and LLMs are provided in Appendix A.1.
@@
The above analysis reveals that Cor-RAIT generates numerous similar sample pairs between $D_{\text{van}}$ and $D_{\text{idk}}$, resulting in conflicting supervision signals which leads to over-refusal. We term this static conflict to distinguish it from another conflict type discussed later.
@@ 3.3 Certainty Mitigates the Static Conflicts
We conducted a theoretical analysis 333detailed proof in Appendix A.2. establishing a weak (non-differentiable) link between the LLM’s feature and the response correctness $\mu$ for the specific question $q$. This weak correlation causes highly similar samples being categorized into $D_{\text{van}}$ and $D_{\text{idk}}$ respectively. To mitigate this, we propose incorporating a robust indicator variable aligned with correctness to select and construct the RAIT data. This variable should ensure that similar samples share comparable values, reducing the above misclassification. We suggest adopting the certainty (Jiang et al. 2023) of the LLM’s response as the indicator variable. Our theoretical analysis shows that certainty meets the above requirements. [^4] Section 4.2 discusses various methods for representing LLM response certainty. Appendix A.3 uses entropy as a measure, but our findings extend to other methods as well.
@@ 3.4 Knowledge Flow and Dynamic Conflict
Research (Ren et al. 2024; Gekhman et al. 2024; Ren and Sutherland 2024) reveals that the knowledge state of LLMs evolves during the SFT process. The phenomenon, which we refer to as “knowledge flow”, can cause previously incorrectly answered questions to become correct ones during SFT. Despite this dynamic evolutions, the target answer of the training data remains static during the RAIt, which reflects the knowledge state of the initial LLM but ignores subsequent changes. We term it as dynamic conflict, which significantly contributes to the over-refusal in Cor-RAIT.
We select the data with the highest correctness and certainty for SFT, a process we refer to as rehearsal training [^5] Details are provided in Appendix B.2.. Rehearsal training is designed to capture the LLM’s natural knowledge flow during SFT. Experiments on the MMLU dataset (Hendrycks et al. 2020) and LLaMA3-8B-Instruct demonstrated that the correctness of 69% of samples, initially below 0.5, improved, thereby validating the aforementioned analysis on dynamic conflict. Additional experimental results are provided in Appendix A.5.